# Online Fair Allocations with Binary Valuations and Beyond

### v1

@@ 5 Conclusion and Future Work
In our paper, we study the fair online allocation of indivisible items and present a series of possible and impossible results regarding the existence and approximation results in fairness and efficiency with various valuation or cost functions, which give a picture of what can be achieved. For future work, one interesting direction is to fill the gap of the approximation of EF1 and MMS for bi-valued valuation or cost functions when there are more than three agents. Additionally, based on our results in other online settings, it would also be interesting to study the items with deadline setting in a more complicated valuation or cost function setting.
Appendix A Missing Proofs
A.1 Section 3
A.1.1 Proof of Proposition 5
Proposition 5.
For general additive functions, no deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any $\alpha>0$ when there are more than two agents with additive tri-valued valuation functions.
Proof.
$t=1$ $t=2$ldots $t=n-2$ $t=n-1$ $t=n$ $t=n+1$
agent $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent $2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
agent $n-2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent $n-1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
More than two agents
Table 10: The impossibility results of EF1 and MMS for additive tri-valued valuations.
We further extend the impossibility results of approximate EF1 and MMS for $n\geq 3$. Consider the following instance with $n\geq 3$ agents, where the value of each item is shown in Table 10. In the first $(n-2)$ rounds, to avoid 0-EF1 or 0-MMS, each item is allocated to the agent with the empty bundle. For the item that arrives before round $t=n-2$, it has the value of $\frac{1}{\epsilon}$ for any agent. Without loss of generality, we assume that agent $i$ picks the item in the round $t=i$, where $i\leq n$. In round $t=n-1$, the arriving item has the value of 1 for any agent, and we assume that it is allocated to agent $n-1$. In round $t=n$, the arriving item has the value of $\epsilon$ for agent $i\in N\setminus\{n-1\}$ and $\frac{1}{\epsilon}$ for agent $n-1$, and it should be allocated to agent $n$ to avoid 0-EF1 or 0-MMS. Finally, in the round $t=n+1$, the item with the value of $\frac{1}{\epsilon}$ for each agent arrives. Due to non-wastefulness, this item must be allocated to one agent. If this item is allocated to agent $i\in N\setminus\{n-1,n\}$, it is not hard to see that the allocation is $\epsilon^{2}$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS. If this item is allocated to agent $n-1$, this allocation is $\epsilon$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS. If this item is allocated to agent $n$, we can derive that the competitive ratio of EF1 is $\epsilon$ and MMS is $\epsilon$. Combining the above cases together, we can conclude that the upper bound of EF1 is $\epsilon$ and MMS is $\epsilon$. For given $\alpha>0$, let $\epsilon<\alpha$, we can conclude that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS for any given $\alpha>0$. ∎
A.2 Section 4
A.2.1 Proposition 6
Proposition 6.
For general additive cost functions, no deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$ for more than two agents with additive tri-valued cost functions.
Proof.
$t=1$ $t=2$ldots $t=n-2$ $t=n-1$ $t=n$ $t=n+1$
agent $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent $2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
agent $n-2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent $n-1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
More than two agents
Table 11: The impossibility result of EF1 for additive tri-valued cost functions.
We extend the impossibility result of approximate EF1 for chores to $n\geq 3$ agents. Consider the following instance with $n\geq 3$ agents, where the cost of each item is shown in Table 11. Before each agent’s bundle has one agent, we should allocate the arriving item to the agent with an empty bundle to avoid the approximation ratio of EF1 being infinite. For $t\leq n-2$, each arriving item has the cost of $\frac{1}{\epsilon}$ for each agent. Without loss of generality, assume that agent $i\in[n-2]$ picks the item that arrives at round $i$. When $t=n-1$, the arriving item has the cost of 1 for all agents. Assume that it is allocated to agent $n-1$. When $t=n$, the arriving item has the cost of $\frac{1}{\epsilon}$ for agent $n$ and $\epsilon$ for the remaining agents. It is clear that agent $n$ picks it. In the last round $t=n+1$, the item that has the cost of $\frac{1}{\epsilon}$ for all agents arrives. If it is allocated to agent $i\in[n-2]$, we can derive that the approximation ratio of EF1 is $\frac{1}{\epsilon}$. If it is allocated to agent $n-1$ or $n$, the allocation is also $\frac{1}{\epsilon}$-EF1. Thus, the allocation in this round is $\frac{1}{\epsilon}$-EF1. Given any $\alpha>1$, let $\epsilon>\frac{1}{\alpha}$, we can conclude that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha>1$. ∎
Appendix B Other Online Settings
@@
Online algorithms face two roadblocks: unknown future information and the inability to revoke previous decisions. A natural question is how to avoid the worst-case scenario in online algorithms. In this section, we study two special settings. One is that each arriving item can wait for one period and does not need to be allocated immediately, which can be found in the online matching literature [23]. In this setting, we prove that an EF1 allocation for two agents with bi-valued valuation/cost functions exists. The other one is knowing some information in advance, which originates from the online MMS allocation with normalized valuations [5]. In this setting, we show that the Marginal-Greedy Algorithm computes a non-wasteful EF1 allocation when items are goods, and the Round-Robin algorithm computes a complete EF1 allocation when items are chores.
B.1 Items with Deadline
@@
In this part, each item has a deadline $d$ after its arrival. After $d$ units of time, we must decide how to allocate it. For two agents with bi-valued valuation functions in the goods setting, we design the Adapted Matching Algorithm (Algorithm 12) to compute an NW allocation that satisfies EF1. For two agents with bi-valued cost functions in the chores setting, we design the Adapted Chores Matching Algorithm (Algorithm 13) to compute a complete allocation that satisfies EF1. The common intuition of the above two algorithms is that we find a perfect matching between two agents and two items after two consecutive items arrive so that we can guarantee that there is no envy cycle between two agents and EF1 always holds.
@@
For the deterministic allocation of indivisible goods for two agents with bi-valued valuation functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Matching Algorithm (Algorithm 12) computes an NW allocation that satisfies EF1.
@@
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 12. We will show that the returned allocation $A$ is always EF1 by induction. For the base case, the empty bundle is trivially EF1 without an envy cycle. For the induction step, after allocating item $e_{t}$ for any $t\in[k]$, the allocation $A^{t}$ is EF1 without an envy cycle. Next, we show that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is EF1 without an envy cycle. Let us consider three cases about item $e_{k+1}$.
  • $\lambda=1$ and item $e_{k+1}$ is the last arriving item. When envy exists between two agents, assume that agent $i$ is the agent who envies agent $j=2-i$. If she picks $e_{k+1}$, we are done. If not, agent $j$ picks it and we get $v_{i}(e_{k+1})=a_{i}$. For agent $j$, she still does not envy agent $i$, and her new bundle is EF. For agent $i$, by Lemma 6, we have $v_{i}(A_{i}^{k})-v_{i}(A_{j}^{k})=a_{i}-b_{i}$. Then, we have $v_{i}(A_{i}^{k})-v_{i}(A_{j}^{k}\cup\{e_{k+1}\}\setminus\{g\})=a_{i}-b_{i}-a_{i}+b_{i}=0$ where $A_{i}^{k}$ is EF1 and there must exists an item $g\in A_{j}$ such that $v_{i}(g)=b_{i}$, implying the allocation $A^{k+1}$ is EF1 without an envy cycle. When no envy exists between two agents, allocating $e_{k+1}$ to any agent will not violate EF1.
  • $\lambda=2$ and item $e_{k+1}$ needs to be allocated firstly. When envy exists between two agents, assume that agent $i$ is the agent who envies agent $j=2-i$. If agent $i$ picks $e_{k+1}$, we are done. If not, for agent $i$, her bundle is still EF1 since we have $v_{i}(A_{i})-v_{i}(A_{j}\cup\{e_{k+1}\}\setminus\{g\})=a_{i}-b_{i}-a_{i}+b_{i}=0$ where there must exist $g\in A_{j}$ such that $v_{i}(g)=b_{i}$. For agent $j$, she does not envy agent $i$ and by Lemma 6, we have $v_{j}(A_{j})-v_{j}(A_{i})\geq b_{j}-a_{j}$. Thus, her bundle is EF. When no envy exists between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
  • $\lambda=2$ and item $e_{k+1}$ needs to be allocated secondly. When envy exists between two agents, assume that agent $i$ is the agent who envies agent $j=2-i$. If she picks it, we are done. If not, it means that item $e_{k}$ is allocated to agent $i$ and $v_{i}(e_{k})\geq v_{i}(e_{k+1})$. For agent $i$, considering that $A_{i}^{k-1}$ is EF1, we can derive that her bundle is still EF1. For agent $j$, it is trivial that her new bundle is EF. When no envy exists between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
Combining the above cases together, we can conclude that the allocation $A^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
@@
For the deterministic allocation of indivisible chores for two agents with bi-valued cost functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Chores Matching algorithm (Algorithm 13) computes a complete allocation that satisfies EF1.
@@
  • $c_{i}(A_{i})<c_{i}(A_{j})$. After allocating two items $e_{k+1}$ and $e_{k+2}$, the two properties hold since agent $i$ only picks one item and has a bi-valued cost function i.e., $c_{i}(e_{k+1}),c_{i}(e_{k+2})\in\{a_{i},b_{i}\}$.
@@
Given an online fair allocation instance $I=(T,N,(c_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 13. We will show that the output allocation $A$ is always EF1 by induction. For the base cases, the empty bundle is EF1 without an envy cycle. For the induction step, after allocating item $e_{t}$ for any $t\in[k]$, the allocation $A^{t}$ is EF1 without an envy cycle. Next, we show that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is EF1 without an envy cycle. Let us consider three cases about item $e_{k+1}$.
@@
  • $\lambda=1$ and item $e_{k+1}$ is the last arriving item. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent $j=2-i$. If she picks it, we are done. If agent $j$ picks it, it means that $c_{j}(e_{k+1})=a_{j}$. For agent $i$, she does not envy agent $j$, so her bundle is still EF. For agent $j$, by Lemma 7, we have $c_{j}(A_{j}^{k})-c_{j}(A_{i}^{k})=b_{j}-a_{j}$. Then, we have $c_{j}(A_{j}^{k}\setminus\{f\}\cup\{e_{k+1}\})-c_{j}(A_{i})=b_{j}-a_{j}-b_{j}+a_{j}=0$ where $A_{j}^{k}$ is EF1 and there must exist an item $f$ such that $c_{j}(f)=b_{j}$, implying that allocation $A^{k+1}$ is EF1 without an envy cycle. When there is no envy between two agents, allocating $e_{k+1}$ to any agent will not violate EF1.
@@
  • $\lambda=2$ and item $e_{k+1}$ needs to allocated firstly. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent $j=2-i$. If agent $i$ picks it, we are done. If not, agent $j$ picks it. For agent $i$, she does not envy agent $j$, so her bundle is EF. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}^{k})-c_{j}(A_{i}^{k})=b_{j}-a_{j}$ and $c_{j}(e_{k+1})=a_{j}$ by the condition of our algorithm. Thus, this allocation is EF1 without an envy cycle. When there is no envy between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
@@
  • $\lambda=2$ and item $e_{k+2}$ needs to be allocated secondly. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent $j=2-i$. If agent $i$ picks it, we are done. If not, agent $j$ picks it. In that case, item $e_{k}$ is allocated to agent $i$. For agent $i$, she does not envy agent $j$, so her bundle is EF. For agent $j$, we can still derive that her new bundle is EF1 since by our assumption, we have $A^{k-1}_{j}$ is EF1, and $c_{j}(e_{k+2})\leq c_{j}(e_{k+1})$ by the condition of our algorithm. Thus, this allocation is EF1 without an envy cycle. When there is no envy between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
@@
Combining the above cases together, we can conclude that the allocation $A^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
B.2 Partial Information in Advance
Definition 12.
An online goods/chores instance is monotone, if for any agent $i\in N$, we have $f_{i}(e_{1})\geq\cdots\geq f_{i}(e_{T})$ or for any agent $i\in N$, we have $f_{i}(e_{1})\leq\cdots\leq f_{i}(e_{T})$, where $f_{i}\in\{v_{i},c_{i}\}$
Theorem 17.
Proof.
Without loss of generality, we assume that the picking sequence is $1,\ldots,n$. First, let us consider the case of $v_{i}(e_{1})\leq\ldots\leq v_{i}(e_{T})$ for any $i\in N$. Note that, for any agent $i\in N$, if she values some items at zero, these items will arrive before the items with non-zero values. In this phase, she does not pick anything until the item with non-zero values arrives. Additionally, when these zero items are allocated to other agents, from the perspective of agent $i$, the value of the other agent’s bundle remains unchanged. Thus, without loss of generality, we assume that $v_{i}(e_{1})\neq 0$ for any agent $i\in N$. Fix an arbitrary round $k$ and an arbitrary agent $i\in N$. Let $A^{k}$ denote the allocation after the allocation of $e_{k}$. Suppose that she picks item $e_{k}$. For agent $i$, we have $v_{i}(A^{k}_{i})\geq v_{i}(A^{k}_{j}\setminus\{e\})$, where $e$ is the last item added to agent $j$’s bundle, since for any $j<i$, agent $i$ is the latter one to pick the item and for any $j>i$, $v_{i}(e_{s})\geq v_{i}(e_{t})$ holds, where agent $i$ picks $e_{s}$, agent $j$ picks $e_{t}$ and $s-t=n-i+j$. For agent $j\neq i$, we have $v_{j}(A_{j}^{k})\geq v_{j}(A_{i}^{k}\setminus\{e\})$, where $e$ is the last item added to agent $i$’s bundle, since for any $j>i$, $v_{j}(e_{s})\geq v_{j}(e_{t})$, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and $s-t=j-i$, and $j<i$, $v_{j}(e_{s})\geq v_{j}(e_{t})$ holds, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and $s-t=n-j+i$. Thus, the allocation $A^{k}$ is EF1. As for the case of $v_{i}(e_{1})\geq\ldots\geq v_{i}(e_{T})$ for any $i\in N$, we can use a similar analysis to show that the returned allocation is EF1 and we omit it. ∎
Theorem 18.
Proof.
Without loss of generality, we assume that the picking sequence is $1,\ldots,n$. First, let us consider the case of $c_{i}(e_{1})\leq\ldots\leq c_{i}(e_{T})$ for any $i\in N$. Fix an arbitrary round $k$ and an arbitrary agent $i\in N$. Let $A^{k}$ denote the allocation after the allocation of $e_{k}$. Suppose that she picks item $e_{k}$. For agent $i$, we have $c_{i}(A^{k}_{i}\setminus\{e_{k}\})=c_{i}(A^{k-1}_{i})\leq c_{i}(A^{k-1}_{j})=c_{i}(A^{k}_{j})$ since for any $j>i$, agent $i$ has the priority of picking the item and for any $j<i$, $c_{i}(e_{s})\leq c_{i}(e_{t})$ holds, where agent $i$ picks $e_{s}$, agent $j$ picks $e_{t}$ and $t-s=n-i+j$. For agent $j\neq i$, we have $c_{j}(A_{j}^{k})\leq c_{j}(A_{i}^{k})$ for any $j<i$ agent $j$ has the priority of picking items and $j>i$, $c_{j}(e_{s})\leq c_{j}(e_{t})$ holds, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and $t-s=n-j+i$. Therefore, the allocation $A^{k}$ is EF1. As for the case of $c_{i}(e_{1})\geq\ldots\geq c_{i}(e_{T})$ for any $i\in N$, we can use a similar analysis to show that the returned allocation is EF1 and we omit it.∎
Proposition 7.
For an online goods/chores instance where agents only have identical preferences, there exists an instance for which no EF1 allocation can be found, even for two agents with bi-valued valuation or cost functions.
Proof.
$t=1$ $t=2$ $t=3$ $t=4$
agent $1$ $1$ $1$ $\epsilon$ $1$
agent 2 $1$ $\epsilon$ $\epsilon$ $1$
Case 1
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
agent $1$ $1$ $1$ $\epsilon$ $1$ $1$
agent 2 $1$ $\epsilon$ 1 $\epsilon$ $1$
Case 2
$t=1$ $t=2$ $t=3$
agent $1$ $1$ $1$ $\frac{1}{\epsilon}$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 1
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
agent $1$ $1$ $1$ $\frac{1}{\epsilon}$ 1 $\frac{1}{\epsilon}$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 2
First, we consider the online goods instance. Without loss of generality, we assume that $v_{1},v_{2}\in\{\epsilon,1\}$, where $10<\epsilon<$ 1 The value of each item is shown in Table 12. When $t=1$, the arriving item has the value of $1$ for both agents. Without loss of generality, it is allocated to agent 1. For the second round $t=2$, the item that has the value of $1$ for agent 1 and $\epsilon$ for agent 2. To guarantee EF1, agent 2 picks it. In round $t=3$, the arriving item has the value of $\epsilon$ for both agents. Then, let us consider two cases related to allocating it.
Next, let us consider the online chores instance. We assume that $c_{1},c_{2}\in\{1,\frac{1}{\epsilon}\}$ where $10<\epsilon<$ 1. The cost of each item is shown in Table 13. When $t=1$, the arriving item has the cost of 1 for both agents, and we assume that agent 1 picks it. In round $t=2$, the item that has the cost of 1 for agent $1$ and $\frac{1}{\epsilon}$ for agent 2. To maintain EF1, agent 2 picks it. For the next round $t=3$, the arriving item has the cost of $\frac{1}{\epsilon}$ for both agents. Then, let us consider two cases related to allocating it.
Combining the above cases, we can conclude that our statement holds. ∎
@@ B.2 Partial Information in Advance
Input: An instance $(T,N,(v_{1},v_{2}))$ with bi-valued valuation functions and deadline $d=1$
Output: An non-wasteful EF1 allocation $A$

### v2

@@ 5 Conclusion and Future Work
In our paper, we study the fair online allocation of indivisible items and present a series of positive and negative results regarding the existence and approximation results in fairness and efficiency with various valuation or cost functions, which provide a comprehensive view of what can be achieved. For future work, the most interesting direction is to fill the gap in the approximation of EF1 and MMS for bi-valued valuation or cost functions when there are more than two agents, because it involves the complicated envy cycle, and the current technique cannot be directly generalized. Additionally, based on our results in other online settings, it would also be interesting to study the items with deadline setting in a more complicated valuation or cost function setting.
Appendix A Other Online Settings
@@
Online algorithms face two roadblocks: unknown future information and the inability to revoke previous decisions. A natural question is how to avoid the worst-case scenario in online algorithms. In this section, we study two special settings. One is that each arriving item can wait for one period and does not need to be allocated immediately, which can be found in the online matching literature Ashlagi et al. [2018]. In this setting, we prove that an EF1 allocation for two agents with additive personalized bi-valued valuation/cost functions exists. The other one is knowing some information in advance, which originates from the online MMS allocation with normalized valuations Zhou et al. [2023]. In this setting, we show that the Marginal-Greedy Algorithm computes a non-wasteful EF1 allocation when items are goods, and the Round-Robin algorithm computes a complete EF1 allocation when items are chores.
A.1 Items with Deadline
@@
In this part, each item has a deadline $d$ after its arrival. After $d$ units of time, we must decide how to allocate it. For two agents with additive personalized bi-valued valuation functions in the goods setting, we design the Adapted Matching Algorithm (Algorithm 14) to compute an NW allocation that satisfies EF1. For two agents with additive personalized bi-valued cost functions in the chores setting, we design the Adapted Chores Matching Algorithm (Algorithm 15) to compute a complete allocation that satisfies EF1. The common intuition of the above two algorithms is that we find a perfect matching between two agents and two items after two consecutive items arrive so that we can guarantee that there is no envy cycle between two agents and EF1 always holds.
@@
For the deterministic allocation of indivisible goods for two agents with additive personalized bi-valued valuation functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Matching Algorithm (Algorithm 14) computes an NW allocation that satisfies EF1.
@@
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$is the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 14.
EF1.
We will show that the returned allocation A is always EF1 by induction. For the base case, the empty bundle is trivially EF1 without an envy cycle. For the induction step, after allocating item $e_{t}$ for any $t\in[k]$, the allocation $A^{t}$ is EF1 without an envy cycle. Next, we show that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is EF1 without an envy cycle. Let us consider three cases about item $e_{k+1}$.
  • $\lambda=1$ and item $e_{k+1}$ is the last arriving item. When envy exists between two agents, assume that agent $i$ is the agent who envies agent j=2ij=2-i. If she picks $e_{k+1}$, we are done. If not, agent $j$ picks it and we get $v_{i}(e_{k+1})=a_{i}$. For agent $j$, she still does not envy agent $i$, and her new bundle is EF. For agent $i$, by Lemma 11, we have $v_{i}(A_{i}^{k})-v_{i}(A_{j}^{k})=a_{i}-b_{i}$. Then, we have $v_{i}(A_{i}^{k})-v_{i}(A_{j}^{k}\cup\{e_{k+1}\}\setminus\{g\})=a_{i}-b_{i}-a_{i}+b_{i}=0$ where $A_{i}^{k}$ is EF1 and there must exists an item $g\in A_{j}$ such that $v_{i}(g)=b_{i}$, implying the allocation $A^{k+1}$ is EF1 without an envy cycle. When no envy exists between two agents, allocating $e_{k+1}$ to any agent will not violate EF1.
  • $\lambda=2$ and item $e_{k+1}$ needs to be allocated firstly. When envy exists between two agents, assume that agent $i$ is the agent who envies agent j=2ij=2-i. If agent $i$ picks $e_{k+1}$, we are done. If not, for agent $i$, her bundle is still EF1 since we have $v_{i}(A_{i})-v_{i}(A_{j}\cup\{e_{k+1}\}\setminus\{g\})=a_{i}-b_{i}-a_{i}+b_{i}=0$ where there must exist $g\in A_{j}$ such that $v_{i}(g)=b_{i}$. For agent $j$, she does not envy agent $i$ and by Lemma 11, we have $v_{j}(A_{j})-v_{j}(A_{i})\geq b_{j}-a_{j}$. Thus, her bundle is EF. When no envy exists between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
  • $\lambda=2$ and item $e_{k+1}$ needs to be allocated secondly. When envy exists between two agents, assume that agent $i$ is the agent who envies agent j=2ij=2-i. If she picks it, we are done. If not, it means that item $e_{k}$ is allocated to agent $i$ and $v_{i}(e_{k})\geq v_{i}(e_{k+1})$. For agent $i$, considering that $A_{i}^{k-1}$ is EF1, we can derive that her bundle is still EF1. For agent $j$, it is trivial that her new bundle is EF. When no envy exists between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
Combining the above case, we can conclude that the allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
@@
For the deterministic allocation of indivisible chores for two agents with additive personalized bi-valued cost functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Chores Matching algorithm (Algorithm 15) computes a complete allocation that satisfies EF1.
@@
  • $c_{i}(A_{i})<c_{i}(A_{j})$. After allocating two items $e_{k+1}$ and $e_{k+2}$, the two properties hold since agent $i$ only picks one item and has a additive personalized bi-valued cost function i.e., $c_{i}(e_{k+1}),c_{i}(e_{k+2})\in\{a_{i},b_{i}\}$.
@@
Given an online fair allocation instance $I=(T,N,(c_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 15.
EF1.
@@
We will show that the output allocation A is always EF1 by induction. For the base cases, the empty bundle is EF1 without an envy cycle. For the induction step, after allocating item $e_{t}$ for any $t\in[k]$, the allocation $\textbf{A}^{t}$ is EF1 without an envy cycle. Next, we show that after allocating item $e_{k+1}$, the allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle. Let us consider three cases about item $e_{k+1}$.
@@
  • $\lambda=1$ and item $e_{k+1}$ is the last arriving item. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent j=2ij=2-i. If she picks it, we are done. If agent $j$ picks it, it means that $c_{j}(e_{k+1})=a_{j}$. For agent $i$, she does not envy agent $j$, so her bundle is still EF. For agent $j$, by Lemma 12, we have $c_{j}(A_{j}^{k})-c_{j}(A_{i}^{k})=b_{j}-a_{j}$. Then, we have $c_{j}(A_{j}^{k}\setminus\{f\}\cup\{e_{k+1}\})-c_{j}(A_{i})=b_{j}-a_{j}-b_{j}+a_{j}=0$ where $A_{j}^{k}$ is EF1 and there must exist an item $f$ such that $c_{j}(f)=b_{j}$, implying that allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle. When there is no envy between two agents, allocating $e_{k+1}$ to any agent will not violate EF1.
@@
  • $\lambda=2$ and item $e_{k+1}$ needs to allocated firstly. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent j=2ij=2-i. If agent $i$ picks it, we are done. If not, agent $j$ picks it. For agent $i$, she does not envy agent $j$, so her bundle is EF. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}^{k})-c_{j}(A_{i}^{k})=b_{j}-a_{j}$ and $c_{j}(e_{k+1})=a_{j}$ by the condition of our algorithm. Thus, this allocation is EF1 without an envy cycle. When there is no envy between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
@@
  • $\lambda=2$ and item $e_{k+2}$ needs to be allocated secondly. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent j=2ij=2-i. If agent $i$ picks it, we are done. If not, agent $j$ picks it. In that case, item $e_{k}$ is allocated to agent $i$. For agent $i$, she does not envy agent $j$, so her bundle is EF. For agent $j$, we can still derive that her new bundle is EF1 since by our assumption, we have $A^{k-1}_{j}$ is EF1, and $c_{j}(e_{k+2})\leq c_{j}(e_{k+1})$ by the condition of our algorithm. Thus, this allocation is EF1 without an envy cycle. When there is no envy between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
@@
Combining the above cases, we can conclude that the allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
@@
Input: An instance $(T,N,(v_{1},v_{2}))$ with additive personalized bi-valued valuation functions and deadline $d=1$
Output: An NW EF1 allocation A